OpenStudy (shaleiah):

Check my answer? @Hicklyfe123

OpenStudy (shaleiah):

A construction company orders 2900 new light bulbs for a new office building. If the probability of a light bulb being defective is 1.9%, predict how many of the light bulbs are likely to be defective. Round your answer to the nearest whole number.

OpenStudy (shaleiah):

55 light bulbs c. 46 light bulbs b. 69 light bulbs d. 551 light bulbs

OpenStudy (misty1212):

HI!!

OpenStudy (shaleiah):

1.9/100*2900=55

OpenStudy (shaleiah):

Hello

OpenStudy (misty1212):

muliply \[0.019\times 2900

OpenStudy (misty1212):

oops \[0.019\times 2900\]

OpenStudy (misty1212):

yeah 55

OpenStudy (hicklyfe123):

okay do u know pemdas?

OpenStudy (misty1212):

yeah dated his brother

OpenStudy (hicklyfe123):

Oh funny.... i was asking shaleiah

OpenStudy (hicklyfe123):

@shaleiah your correct PEMDAS... 1.9/100*2900= your answer :)

OpenStudy (shaleiah):

Sorry my computer froze :c

OpenStudy (shaleiah):

Thank you both!